OpenStudy (anonymous):

Factor x^3-x^2-x+1 I got the first part I think but don't know what to do with the -x X^2(x-1) ? @ precal

OpenStudy (anonymous):

use factor theorem: if f(a) = 0 then (x-a) is a factor f(1) = 0 hence (x-1) is a factor (x-1)(ax^2 + bx + c) we can see that a = 1 ( to get x^3) and c = -1 (to get 1) (x-1)(x^2 + bx -1) = x^3-x^2-x+1 compare coefficients of x: -b - 1 = -1 therefore b = 0 now difference of two squares applies to what we have left: (x^2 -1)(x-1) = (x+1)(x-1)(x-1) = (x+1)(x-1)^2